# Cost-sensitive computational adequacy of higher-order recursion in synthetic domain theory

### v1

@@ Abstract
Acknowledgements. This work was funded by the United States Air Force Office of Scientific Research under grants FA9550-23-1-0728111Dr. Tristan Nguyen, Program Manager (New Spaces for Denotational Semantics), MURI FA9550-21-0009Footnote 1, FA9550-21-1-0385Footnote 1 and the National Science Foundation under grant CCF-1901381. We thank Tristan Nguyen at AFOSR for support. Yue Niu was supported by the Air Force Research Laboratory through the NDSEG fellowship. Views and opinions expressed are however those of the author only and do not necessarily reflect those of AFOSR, AFRL, or NSF.
@@ 1.2.1 Cost as a phase distinction
From an external point of view, calf’s types can be modeled as presheaves on the interval ${\left\{\mathsf{ext}\sqsubseteq\mathsf{int}\right\}}$; in the terminology of Strling and Harper [30], such a presheaf encodes a set varying in the intension-extension “security poset” ${\left\{\mathsf{ext}\sqsubseteq\mathsf{int}\right\}}$ whose presheaf restriction action redacts intensional cost structure.
@@ 1.2.2 Synthetic domain theory
Synthetic domain theory (SDT) as a field started when Dana Scott conjectured that one ought to be able to reason about domains as if they were just sets provided that one employs a constructive ambient metalanguage. In technical terms, this metalanguage can be construed as the internal languages of a topos, i.e. an extensional dependent type theory. In general it is difficult to reconcile the domain-theoretic structure with the rich logical structure of dependent type theory, so the quest for SDT was in essence about constructing full subcategories of topoi that supported domain-theoretic constructions. Fullness is a critical property — it means that every map definable in type-theoretic language is a domain morphism, which absolves one from checking onerous side conditions when working internally. Today there are two kinds of synthetic domain theory that are mutually incompatible: one based on realizability (Hyland [10], Phoa [18], Reus [22]) and one based on sheaf topoi (Rosolini [25], Fiore and Rosolini [6], Fiore and Plotkin [4], Matache et al. [15]). In this paper we will construct a sheaf model of SDT, so we take a moment to recall the basic ideas, mainly drawing from the work of Fiore and Plotkin [4].
@@ 1.2.3 Predomains and the intrinsic order
Aside from intrinsic interest, we are motivated by two practical incentives for considering and developing the theory of the synthetic preorder. First, we would like to connect our work to the type theory for cost analysis developed by Grodin et al. [7]. In that setting one works with a cost ordering on programs in which $e\sqsubseteq e^{\prime}$ represents a proof of the fact that $e^{\prime}$ is an upper bound of $e$ with respect to computational cost. We expect that one may extend the work of op. cit. to account for recursive programs by the methods we develop in this paper, and the information/intrinsic preorder appears to be the correct theoretical framework for describing the interaction of cost and partiality [12, Section 8]. Second, the advantages of the synthetic point of view outlined Section 1.2.2 do not detract from the benefits of the language of classical domain theory. The intrinsic preorder integrates the advantages of both the synthetic and classical perspective by providing an intuitively appealing language for reasoning about domains that automatically complies with well-definedness conditions such as monotonicity and continuity.
@@ 2 A type theory for cost-sensitive synthetic domain theory
By a domain we shall mean a predomain $X$ equipped with a least element. Equivalently this may be characterized as an algebra for the lifting monad $\mathbb{L}=(\mathsf{L},\eta,\mu)$, which freely adjoins a least element to a predomain. There is a special domain $\Sigma\cong\mathsf{L}1\hookrightarrow\Omega$ spanned by propositions called the Sierpiński space or dominance; propositions in $\Sigma$ should be thought of as domains of definition of partial maps, and a map $X\to\Sigma$ can be thought of as a “computational” or “observable” subset of a predomain $X$. Moreover, because we use the sealing modality (see Section 1.2.1) associated the the phase proposition $\P$ to define the denotational cost semantics of PCF, we require that predomains are closed under the sealing modality. Lastly, we assume that the subuniverse $\mathcal{U}_{\mathsf{predom}}$ is closed under lifting.
@@ 3.1 Partial maps, dominance, and lifting
Given a type $A$, the lift of $A$ is defined as $\mathsf{L}A=\Sigma_{\phi:\Sigma}.~{}\phi\to A$.
@@
Given a partial element $e:\mathsf{L}A$, we write $e{\downarrow}:\Sigma$ for its support. When it is known that $e{\downarrow}$ holds, we may write $e:A$ for the defined element.
@@ 3.2 Complete types
Complete types are the synthetic counterpart to $\omega$-cpos in classic domain theory. The class of well complete types is the least restrictive possible notion of predomain is closed under lifting. In this paper we consider the dual most restrictive class of predomain, the replete types, in order to obtain a sharper characterisation of the intrinsic order relation (see Section 3.3) on predomains.
@@ 3.4 Axioms of the phase distinction in SDT and predomains
[…]
@@ 4.1 Discrete predomains
A type $A$ is called discrete when $x\sqsubseteq y$ implies $x=y$.
@@
A type is called $\Sigma$-decidable when its equality relation is valued in $\Sigma$-propositions.
@@
Any $\Sigma$-decidable type $A$ is discrete.
@@
The category of predomains possesses a natural numbers type ${\mathbb{N}}_{\mathsf{P}}$ that is $\Sigma$-decidable and thus discrete. Note that it is not necessarily the same as the ambient natural numbers type ${\mathbb{N}}$, and we will not assume that it is the case in our constructions. From a logical perspective, this difference means that ${\mathbb{N}}_{\mathsf{P}}$ has a universal mapping out property whose motive is valued in predomains rather than arbitrary types.
@@
If $A$ is a $\Sigma$-decidable type, then so is $\P\vee A$.
@@ 5 $\textbf{PCF}_{\mathsf{cost}}$: a language for cost-aware higher-order recursion
Our main technical result is the computational adequacy property for $\textbf{PCF}_{\mathsf{cost}}$, a version of Plotkin’s PCF [20] equipped with an abstract cost effect. The treatment of both cost structure and recursion as a call-by-push-value effect in $\textbf{PCF}_{\mathsf{cost}}$ is inspired by [12]. The syntax of $\textbf{PCF}_{\mathsf{cost}}$ is parameterized in a monoid $\mathbb{C}$ representing the cost structure; in Section 6 we will impose further properties on $\mathbb{C}$ when we define the denotational semantics of $\textbf{PCF}_{\mathsf{cost}}$. As mentioned in Section 2.1, the type structure of $\textbf{PCF}_{\mathsf{cost}}$ is generated by a pair of operators $\mathsf{F},\mathsf{U}$ that corresponds semantically to the free-forgetful adjunction between plain sets and the category of partial cost algebras, which are sets equipped with an action for the partial cost monad $\mathsf{L}(\mathbb{C}\times-)$. The types and terms of $\textbf{PCF}_{\mathsf{cost}}$ are summarized in Fig. 1.
  • $\textbf{inductive}~{}\mathsf{Ty}^{+}:\mathsf{tp}^{+}~{}\textbf{where}$
@@
  • $\textbf{inductive}~{}\mathsf{Tm}^{+}:\mathsf{Con}\to\mathsf{Ty}^{+}\to\mathsf{tp}^{+}~{}\textbf{where}$
@@
    • $\mathsf{lam}:\mathsf{Tm}^{\ominus}(\mathsf{U}(X)::\Gamma,X)\to\mathsf{Tm}^{\ominus}(\Gamma,X)$
@@ 6 Cost-aware denotational semantics of $\textbf{PCF}_{\mathsf{cost}}$
    Computational: $\mathbb{C}$ is a $\Sigma$-decidable predomain.
@@
[…]
    Discretely ordered: $\mathbb{C}$ is discrete in the sense of Definition 4.2.
@@ 6.2 The derived cost algebra
To model the cost effect $\mathsf{step}:\mathbb{C}\times X\to X$, we use the fact that every algebra for a composite monad is canonically an algebra for the constituent monad. In our case this means that every $\mathbb{T}$-algebra has an underlying cost algebra as well; we write :×$\boxplus:\mathbb{C}\times X\to X$ for the cost algebra map.
@@ 7 Computational semantics of $\textbf{PCF}_{\mathsf{cost}}$
Computational semantics. We consider an alternative formulation of the dynamic semantics of $\textbf{PCF}_{\mathsf{cost}}$ in which execution is modeled not as an inductive family but directly as a (partial) computation. This departure from traditional operational semantics is necessitated by our proof of computational adequacy. As we explain in Section 8.2, termination must be a $\Sigma$-decidable predicate; however, the reflexive-transitive closure of a decidable relation (such as $\mapsto_{A}$) can only be seen to be a $\Sigma$-predicate assuming that $\Sigma$ is closed under countable joins of decidable propositions that are preserved by the inclusion into the ambient type theory, i.e. $\exists[n:{\mathbb{N}}]~{}\phi(n)\in\Sigma$ for every countable family of $\Sigma$-propositions $\phi$. Our approach bypasses this question by de-emphasizing ${\mathbb{N}}$ in favor of the initial lifting algebra $\omega$, a lesson of Simpson [26]. We discuss the relationship between this computational semantics and ordinary operational semantics in Section 11.
@@ 7.1 Computational semantics of $\textbf{PCF}_{\mathsf{cost}}$
We begin with a family of small-step transition relations ${\mapsto_{A}}\subseteq\mathsf{Tm}^{+}(A)\times\mathbb{C}\times\mathsf{Tm}^{+}(A)$ that implements the cost effect model in the sense of Hoffmann [9]. The intuitive meaning of $e\mapsto c,e^{\prime}$ is that $e$ transitions in one step to $e^{\prime}$ and incurs cost $c$; the only place where cost is effected is at $\mathsf{step}$: $\mathsf{step}^{c}(e)\mapsto c,e$. The rest of the rules defining this relation are the standard rules for a call-by-push-value language; we defer to Levy [13] for a detailed account. We iterate the “one step” relation to obtain the partial map implementing the computational semantics. Because $\mapsto_{A}$ is decidable, we have a characteristic map $\mathsf{out}:\{A:\mathsf{tp}^{+}\}.~{}\mathsf{Tm}^{+}(A)\to 1+(\mathbb{C}\times\mathsf{Tm}^{+}(A))$. Consider the following functional:
@@
The relation ${\downarrow}(-)\circ\mathsf{eval}$ is functional.
@@ 8.2 Adequacy
Any $e:\mathsf{U}\mathsf{F}1\to\mathsf{U}\mathsf{F}2$ is weakly, extensionally constant in the sense that for all $x,y:\mathsf{U}\mathsf{F}1$, if $\mathsf{profile}(x){\downarrow}$ and $\mathsf{profile}(y){\downarrow}$, then $\mathsf{eval}(e~{}x,\mathsf{ret}(v)){\downarrow}$ and $\mathsf{eval}(e~{}y,\mathsf{ret}(u)){\downarrow}$ imply $v=u$.
@@ 9 An SDT model of the intension-extension phase distinction
To obtain a model for the constructions of the preceding sections, we instantiate the sheaf model of Sterling and Harper [30] at the poset $\mathbb{I}={\left\{\mathsf{ext}\sqsubseteq\mathsf{int}\right\}}$ representing the intension-extension security order. More explicitly, this corresponds to defining an internal category $\mathscr{C}$ of small dcpos in the presheaf model of the intension-extension of Niu et al. [17] and embedding $\mathscr{C}$ into a Grothendieck topos $\mathrm{Sh}(\mathscr{C})$, obtained by localizing $\widehat{\mathscr{C}}$ at the finite open cover topology. The purpose of this localization is to ensure that the finite joins of the dominance in $\mathscr{C}$ are preserved by the embedding into $\mathrm{Sh}(\mathscr{C})$. This property was notably used by op. cit. to implement the semantics of termination declassification; here we use the finite join structure of $\Sigma$ to show that $2$ is an extensional predomain. The phase distinction in $\mathrm{Sh}(\mathscr{C})$ is inherited from the ambient presheaf topos, where it is represented by $\mathsf{ext}$.
@@
Letting $y:\mathscr{C}\hookrightarrow\mathrm{Sh}(\mathscr{C})$ be the restriction of the Yoneda embedding onto sheaves, we may define $\Sigma=y(\Omega_{\widehat{\mathbb{I}}})$ and $\P=y(\mathsf{y}_{\mathbb{I}}(\mathsf{ext}))$ so that $(\mathrm{Sh}(\mathscr{C}),\Sigma,\P)$ is an SDT model of the intension-extension phase distinction in the sense of Definition 3.16.

### v2

@@ Abstract
Acknowledgements. This work was funded by the United States Air Force Office of Scientific Research under grants FA9550-23-1-0728111Dr. Tristan Nguyen, Program Manager, MURI FA9550-21-0009Footnote 1, FA9550-21-1-0385Footnote 1 and the National Science Foundation under grant CCF-1901381. We thank Tristan Nguyen at AFOSR for support. Yue Niu was supported by the Air Force Research Laboratory through the NDSEG fellowship. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of AFOSR, AFRL, or NSF.
@@ 1.2.1 Cost as a phase distinction
From an external point of view, calf’s types can be modeled as presheaves on the interval $\mathbb{I}={\left\{\mathsf{ext}\sqsubseteq\mathsf{int}\right\}}$; in the terminology of Sterling and Harper [32], such a presheaf encodes a set varying in the intension-extension “security poset” $\mathbb{I}$ whose presheaf restriction action redacts intensional cost structure. The phase proposition $\P$ is given by the “intermediate” proposition $0\to 1$ corresponding to the representable $\mathsf{y}_{\mathbb{I}}(\mathsf{ext})$.
@@ 1.2.2 Synthetic domain theory
Synthetic domain theory (SDT) as a field started when Dana Scott conjectured that one ought to be able to reason about domains as if they were just sets provided that one employs a constructive ambient metalanguage. In technical terms, this metalanguage can be construed as the internal languages of a topos, i.e. an extensional dependent type theory. In general it is difficult to reconcile the domain-theoretic structure with the rich logical structure of dependent type theory, so the quest for SDT was in essence about constructing full subcategories of topoi that supported domain-theoretic constructions. Fullness is a critical property — it means that every map definable in type-theoretic language is a domain morphism, which absolves one from checking onerous side conditions when working internally. There are two well-known ways to obtain models of synthetic domain theory: one based on realizability (Hyland [11], Phoa [20], Reus [24]) and one based on sheaf topoi (Rosolini [27], Fiore and Rosolini [7], Fiore and Plotkin [5], Matache et al. [16]). In this paper we will construct a sheaf model of SDT, so we take a moment to recall the basic ideas, mainly drawing from the work of Fiore and Plotkin [5].
@@ 1.2.3 Predomains and the intrinsic order
Aside from intrinsic interest, we are motivated by two practical incentives for considering and developing the theory of the synthetic preorder. First, we would like to connect our work to the type theory for cost analysis developed by Grodin et al. [8]. In that setting one works with a cost ordering on programs in which $e\sqsubseteq e^{\prime}$ represents a proof of the fact that $e^{\prime}$ is an upper bound of $e$ with respect to computational cost. We expect that one may extend the work of op. cit. to account for recursive programs by the methods we develop in this paper, and the information/intrinsic preorder appears to be the correct theoretical framework for describing the interaction of cost and partiality [13, Section 8]. Second, the advantages of the synthetic point of view outlined Section 1.2.2 do not detract from the benefits of the language of classical domain theory. The intrinsic preorder integrates the advantages of both the synthetic and classical perspective by providing an intuitively appealing language for reasoning about domains that automatically complies with well-definedness conditions such as monotonicity and continuity.
@@ 2 A type theory for cost-sensitive synthetic domain theory
By a domain we shall mean a predomain $X$ equipped with a least element. Equivalently this may be characterized as an algebra for the lifting monad $\mathbb{L}=(\mathsf{L},\eta,\mu)$, which freely adjoins a least element to a predomain. There is a special domain $\Sigma\cong\mathsf{L}1\hookrightarrow\Omega$ spanned by propositions called the Sierpiński space or dominance; propositions in $\Sigma$ should be thought of as the support of partial maps, and a map $X\to\Sigma$ can be thought of as a “computational” or “observable” subset of a predomain $X$. Moreover, because we use the sealing modality (see Section 1.2.1) associated with the phase proposition $\P$ to define the denotational cost semantics of PCF, we require that predomains are closed under the sealing modality. Lastly, we assume that the subuniverse $\mathcal{U}_{\mathsf{predom}}$ is closed under lifting.
@@ 2.1 Denotational semantics of cost-sensitive PCF
Remark 2.2.
As mentioned in Section 1.1.2, we will define both the syntax and semantics of $\textbf{PCF}_{\mathsf{cost}}$ as objects and functions in the internal type theory of a synthetic domain theory topos. In a more traditional approach to denotational cost semantics, one may instead define a model of $\textbf{PCF}_{\mathsf{cost}}$ in the category $\omega\textbf{CPO}^{\to}$ of families of $\omega$-cpos, which is semantically simpler in comparison to models of synthetic domain theory. In this paper we pursue an approach based on the latter in order to unify two strands of prior work — cost-sensitive programming and verification in dependent type theory [18] and internal denotational semantics [19, 17] — and synthetic domain theory provides the means for smoothly integrating higher-order recursion into dependent type theory.
@@ 3.1 Partial maps, dominance, and lifting
The lift of a type $A$ relative to a dominance $\Sigma$ is defined as $\mathsf{L}A=\Sigma_{\phi:\Sigma}.~{}\phi\to A$.
@@
[…]
@@ 3.2 Complete types
Complete types [11, 20] are the synthetic counterpart to $\omega$-cpos in classic domain theory. The class of well complete types was introduced in Longley and Simpson [15] as the least restrictive possible notion of predomain that is closed under lifting. In this paper we consider the dual most restrictive class of predomain, the replete types [11], in order to obtain a sharper characterisation of the intrinsic order relation (see Section 3.3) on predomains.
@@ 3.4 Axioms of the phase distinction in SDT and predomains
An SDT model of the intension-extension phase distinction is an elementary topos $\mathscr{E}$ equipped with a dominance $\Sigma$ satisfying Phoa’s principle such that $\Sigma$ is complete, and a distinguished $\Sigma$-proposition $\P$ such that the type of booleans $2$ is an extensional predomain.
For the rest of the paper we assume a given SDT model of the intension-extension phase distinction $\mathscr{D}$. All constructions are all carried out in the internal language of $\mathscr{E}$.
@@ 4.1 Discrete predomains
A type $A$ is called flat or discrete when $x\sqsubseteq^{\circ}y$ implies $x=y$.
@@
A type has $\Sigma$-equality when its equality relation is valued in $\Sigma$-propositions.
@@
Any type with $\Sigma$-equality is discrete.
@@
The category of predomains possesses a natural numbers type ${\mathbb{N}}_{\mathsf{P}}$ with $\Sigma$-equality, which means it is also discrete. Note that it is not necessarily the same as the ambient natural numbers type ${\mathbb{N}}$, and we will not assume that it is the case in our constructions. From a logical perspective, this difference means that ${\mathbb{N}}_{\mathsf{P}}$ has a universal mapping-out property whose motive is valued in predomains rather than arbitrary types.
@@
If $A$ has $\Sigma$-equality, then so does $\P\vee A$.
@@ 5 $\textbf{PCF}_{\mathsf{cost}}$: a language for cost-aware higher-order recursion
Our main technical result is the computational adequacy property for $\textbf{PCF}_{\mathsf{cost}}$, a version of Plotkin’s PCF [22] equipped with an abstract cost effect. The treatment of both cost structure and recursion as a call-by-push-value effect in $\textbf{PCF}_{\mathsf{cost}}$ is inspired by [13]. The syntax of $\textbf{PCF}_{\mathsf{cost}}$ is parameterized in a monoid $\mathbb{C}$ representing the cost structure; in Section 6 we will impose further properties on $\mathbb{C}$ when we define the denotational semantics of $\textbf{PCF}_{\mathsf{cost}}$. As mentioned in Section 2.1, the type structure of $\textbf{PCF}_{\mathsf{cost}}$ is generated by a pair of operators $\mathsf{F},\mathsf{U}$ that corresponds semantically to the free-forgetful adjunction between plain sets and the category of partial cost algebras, which are sets equipped with an action for the partial cost monad $\mathsf{L}(\mathbb{C}\times-)$. The types and terms of $\textbf{PCF}_{\mathsf{cost}}$ are summarized in Fig. 1, defined as inductive definitions in the SDT topos $\mathscr{E}$.
  • $\textbf{inductive}~{}\mathsf{Ty}^{+}:\mathsf{Set}~{}\textbf{where}$
@@
  • $\textbf{inductive}~{}\mathsf{Tm}^{+}:\mathsf{Con}\to\mathsf{Ty}^{+}\to\mathsf{Set}~{}\textbf{where}$
@@
    • $\mathsf{lam}:\mathsf{Tm}^{\ominus}(A::\Gamma,X)\to\mathsf{Tm}^{\ominus}(\Gamma,A\rightharpoonup X)$
@@ 6 Cost-aware denotational semantics of $\textbf{PCF}_{\mathsf{cost}}$
    Computational: $\mathbb{C}$ is a predomain with $\Sigma$-equality.
@@ 6.2 The derived cost algebra
To model the cost effect $\mathsf{step}:\mathbb{C}\times X\to X$, we use the fact that every algebra for $\mathbb{T}$ is canonically an algebra for the writer monad $\mathbb{C}\times-$, which is a general property of composite monads defined from a distributive law [1, Section 2]. In our case this means that every $\mathbb{T}$-algebra has an underlying cost algebra as well; we write :×$\boxplus:\mathbb{C}\times X\to X$ for the cost algebra map.
@@ 7 Computational semantics of $\textbf{PCF}_{\mathsf{cost}}$
Computational semantics. We consider an alternative formulation of the dynamic semantics of $\textbf{PCF}_{\mathsf{cost}}$ in which execution is modeled not as an inductive family but directly as a (partial) computation. This departure from traditional operational semantics is necessitated by our proof of computational adequacy. As we explain in Section 8.2, termination must be a predicate valued in $\Sigma$-propositions; however, the reflexive-transitive closure of a decidable relation (such as $\mapsto_{A}$) can only be seen to be a $\Sigma$-predicate assuming that $\Sigma$ is closed under countable joins of decidable propositions that are preserved by the inclusion into the ambient type theory, i.e. $\exists[n:{\mathbb{N}}]~{}\phi(n)\in\Sigma$ for every countable family of $\Sigma$-propositions $\phi$. Our approach bypasses this question by de-emphasizing ${\mathbb{N}}$ in favor of the initial lifting algebra $\omega$, a lesson of Simpson [28]. We discuss the relationship between this computational semantics and ordinary operational semantics in Section 11.
@@ 7.1 Computational semantics of $\textbf{PCF}_{\mathsf{cost}}$
[…]
$\mathsf{bind}$($\mathsf{ret}$(a), f) ↦0, f(a)
[…]
$\mathsf{ap}$($\mathsf{lam}$(e), e_1) ↦0, e(e_1)
[…]
$\mathsf{fix}$(e) ↦0, e($\mathsf{fix}$(e))
[…]
$\mathsf{ifz}$($\mathsf{zero}$, e_0, e_1) ↦0, e_0
[…]
$\mathsf{ifz}$($\mathsf{succ}$(v), e_0, e_1) ↦0, e_1(v)
[…]
$\mathsf{step}$^c(e) ↦c, e
Figure 4: Rules for the small-step transition relation.
We begin with a family of small-step transition relations ${\mapsto_{A}}\subseteq\mathsf{Tm}^{+}(A)\times\mathbb{C}\times\mathsf{Tm}^{+}(A)$ that implements the cost effect model in the sense of Hoffmann [10]. The intuitive meaning of $e\mapsto c,e^{\prime}$ is that $e$ transitions in one step to $e^{\prime}$ and incurs cost $c$; the only place where cost is effected is at $\mathsf{step}$: $\mathsf{step}^{c}(e)\mapsto c,e$. The family of relations $\mapsto_{A}$ is defined as an inductive family whose generators are displayed in Fig. 4 (the expected congruence rules have been omitted). We iterate the “one step” relation to obtain the partial map implementing the computational semantics. Because $\mapsto_{A}$ is decidable, we have a characteristic map $\mathsf{out}:\{A:\mathsf{tp}^{+}\}.~{}\mathsf{Tm}^{+}(A)\to 1+(\mathbb{C}\times\mathsf{Tm}^{+}(A))$. Consider the following functional:
@@
The relation $\pi_{1}\circ\mathsf{eval}$ is functional, i.e. ${\downarrow}{\mathsf{eval}(e,v)}$ and ${\downarrow}{\mathsf{eval}(e,v^{\prime})}$ implies $v=v^{\prime}$.
@@ 8.2 Adequacy
Any $e:\mathsf{U}\mathsf{F}1\rightharpoonup\mathsf{F}2$ is weakly, extensionally constant in the sense that for all $x,y:\mathsf{U}\mathsf{F}1$, if $\mathsf{profile}(x){\downarrow}$ and $\mathsf{profile}(y){\downarrow}$, then $\mathsf{eval}(e~{}x,\mathsf{ret}(v)){\downarrow}$ and $\mathsf{eval}(e~{}y,\mathsf{ret}(u)){\downarrow}$ imply $v=u$.
@@ 9 An SDT model of the intension-extension phase distinction
To obtain a model for the constructions of the preceding sections, we instantiate the sheaf model of Sterling and Harper [32] at the poset $\mathbb{I}={\left\{\mathsf{ext}\sqsubseteq\mathsf{int}\right\}}$ representing the intension-extension security order. The basic idea is to first develop domain theory internal to the presheaf topos $\widehat{\mathbb{I}}$, from which we may obtain an appropriate internal domain-theoretic site that embeds into a sheaf topos model of SDT in the sense of Definition 3.16. The reason to consider internal sites is that we may build into the base category the intension-extension phase distinction that is preserved through the embedding.
In op. cit. the internal domain theory of $\widehat{\mathbb{I}}$ is developed in terms of constructive dcpos following the work de Jong [2]. These internal dcpos are similar to ordinary dcpos; for example, one may use them to give the denotational semantics of PCF [2]. The difference lies in the dominance $\Sigma$ of the category of internal dcpos given by the subobject classifier $\Omega_{\widehat{\mathbb{I}}}$: while a partial element of ordinary domains in Set is either defined or not, a partial element of a domain internal to $\widehat{\mathbb{I}}$ may have the phase proposition $\P$ as its support, where $\P$ is the intermediate proposition in $\Omega_{\widehat{\mathbb{I}}}$. Recalling the interpretation of a map $A\to\Sigma$ as a computational predicate, this also means that predicates can be phase-dependent in the sense of holding only at the extensional phase.
We recall some definitions from Sterling and Harper [32] (SH22).
Definition 9.1.
A Scott-open immersion of a dcpo is any mono $U\rightarrowtail A$ arising from a predicate $A\to\Sigma$.
Definition 9.2.
In a category, a sink on an object $A$ is a set of morphisms into $A$.
Definition 9.3.
In a category with pullbacks, a Cartesian coverage is an assignment of objects $A$ to set of sinks on $A$ that is stable under pullback.
Definition 9.4.
The finite open cover topology is generated by the Cartesian coverage assigning to each dcpo $A$ the set of sinks $\{U_{i}\rightarrowtail A\}_{i}$ on $A$ with every $U_{i}\rightarrowtail A$ a Scott-open immersion and $\bigvee_{i}U_{i}\cong A$.
Our domain-theoretic site is given by an internal category $\mathscr{C}$ of small dcpos in $\widehat{\mathbb{I}}$. We embed $\mathscr{C}$ into a Grothendieck topos $\mathrm{Sh}(\mathscr{C})$, obtained by localizing $\widehat{\mathscr{C}}$ at the finite open cover topology. The purpose of this localization is to ensure that the finite joins of the dominance in $\mathscr{C}$ are preserved by the embedding into $\mathrm{Sh}(\mathscr{C})$. This property was notably used by op. cit. to implement the semantics of termination declassification; here we use the finite join structure of $\Sigma$ to show that $2$ is an extensional predomain. The phase distinction in $\mathrm{Sh}(\mathscr{C})$ is inherited from the ambient presheaf topos, where it is represented by $\mathsf{ext}$.
Theorem 9.5 (SH22, Corollary 90).
Every representable presheaf is well-complete.
Thus we obtain a functor $y:\mathscr{C}\hookrightarrow\mathrm{Sh}(\mathscr{C})$ restricting the Yoneda embedding $\mathsf{y}_{\mathscr{C}}:\mathscr{C}\hookrightarrow\widehat{\mathscr{C}}$ onto sheaves.
Theorem 9.6 (SH22, B.1.4.3).
The representable $\Sigma=y(\Sigma)$ is a dominance in $\mathrm{Sh}(\mathscr{C})$.
Theorem 9.7 (SH22, Corollary 79).
Coproducts in $\mathscr{C}$ are given by unions of families of Scott-opens.
Corollary 9.8.
Finite coproducts of $\mathscr{C}$ are preserved by the embedding into sheaves.
Theorem 9.9 (SH22, Axiom-SDT-1).
The dominance $\Sigma$ has finite joins that are preserved by the inclusion $\Sigma\hookrightarrow\Omega$.
@@
Setting $\Sigma=y(\Omega_{\widehat{\mathbb{I}}})$ and $\P=y(\mathsf{y}_{\mathbb{I}}(\mathsf{ext}))$, we have that $(\mathrm{Sh}(\mathscr{C}),\Sigma,\P)$ is an SDT model of the intension-extension phase distinction in the sense of Definition 3.16.
Proof 9.11.
By Theorem 9.5 we know that $\Sigma$ is a well-complete dominance in $\mathrm{Sh}(\mathscr{C})$. To show that $\mathrm{Sh}(\mathscr{C})$ also models the intension-extension phase distinction, we observe that the presheaf model of the intension-extension phase distinction of Niu et al. [18] restricts to a smaller model in $\mathscr{C}$: every subterminal object in $\widehat{\mathbb{I}}$ is an internal dcpo, and so we may take the subterminal $\P=\mathsf{y}_{\mathbb{I}}(\mathsf{ext})$ to be the phase proposition in $\mathscr{C}$. The phase proposition $\P=y(\P)$ in $\mathrm{Sh}(\mathscr{C})$ is classified by $\Sigma$: since $y$ is fully faithful, every $\phi:\Sigma$ arise from a unique map $1_{\widehat{\mathbb{I}}}\to\Omega_{\widehat{\mathbb{I}}}$.
Moreover, we can directly verify that $2_{\mathscr{C}}=2_{\widehat{\mathbb{I}}}=1_{\widehat{\mathbb{I}}}+1_{\widehat{1}}$ is internally orthogonal to $\P$ in $\mathscr{C}$. By Corollary 9.8, $y$ preserves finite coproducts, so we have that $2_{\mathrm{Sh}(\mathscr{C})}=y(2_{\mathscr{C}})$. We observe that 2 is extensional because the restricted embedding $y:\mathscr{C}\to\mathrm{Sh}(\mathscr{C})$ is both full and faithful and preserves products. To see that $2$ is replete, we observe that 2 is isomorphic to its type of singletons:
Because $\Sigma$ is closed under finite joins (by Theorem 9.9), the type of singletons of 2 can be defined as the limit of a diagram of replete types, and so it is replete as well.